OpenStudy (darkigloo):

integral from 0 to 3 abs(x-1) dx

OpenStudy (anonymous):

break it in to two parts, or else use geometry

OpenStudy (anonymous):

geometry is probably easiest

OpenStudy (kainui):

First, to get rid of the absolute value signs I'd solve for when it =0 and then you can make one part that's less than 0 negative to that point and then leave the positive part untouched and adjust your limits of integration accordingly.

OpenStudy (anonymous):



OpenStudy (anonymous):

you have two triangles, left one has base 1 and height 1, right one has base 2 and height 2 use one half base times height to get the area, you can almost do it in your head

OpenStudy (darkigloo):

ok i understand that but can you explain how i would solve it by breaking it into two parts?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x) = |x-1| = \left\{\begin{array}{rcc} 1-x & \text{if} & x <1 \\ x-1& \text{if} & x \geq 1 \end{array} \right. \]

OpenStudy (anonymous):

so \[\int_0^3 f(x)=\int_0^1(1-x)dx+\int_1^3(x-1)dx\]

OpenStudy (anonymous):

i.e. it is a piecewise function, so you have to break the integral in to two pieces this is a rather long way to do it, but it certainly works

OpenStudy (anonymous):

that is why absolute value is such a pain seems like it is easy when you first learn it means "make it positive" but actually since it is a piecewise function it is quite annoying

OpenStudy (anonymous):

much easier to say \(\frac{1}{2}+2=\frac{5}{2}\) and be done with it

OpenStudy (darkigloo):

Ahh i see. Thank you!

OpenStudy (anonymous):

yw